OpenStudy (anonymous):

can u plz help me with some math im in 8th anyone above

OpenStudy (anonymous):

@Umangiasd

OpenStudy (anonymous):

there it is srry bout that

Nnesha (nnesha):

rate of change mean ???

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

can yall tell me wat the slope is for every single one and then i can choose which one is greater

OpenStudy (anonymous):

The rate of change is how much does y vary with x So we can write it as dy/dx now, as they're lines, the rate of change is equal to the slope so; \[\dfrac {dy} {dx} = m=\dfrac {y_2 - y_1} {x_2 - x_1} \] now, you give yourself two points and try out, for example, for: \[x_2=1; x_2 =0\] in equation A, you'll have \[(6*1)-7y_2=21\] \[6-21=7y_2\] \[y_2 = \dfrac {-15} {7}\]

Nnesha (nnesha):

for first equation you all have to just solve for y  subtract 6x both side

OpenStudy (anonymous):

\[(6*0)-7y_1=21\] \[y_1= -3\] So, you have two points, point "2" (1,-15/7) and point "1" (0, -3)

Nnesha (nnesha):

 let me know what you get after this @Darknight200